OpenStudy (anonymous):

Dosage for a drug is 15 mg for a person whose blood volume is 6.0L. (the density of blood is 1.05 g/mL). 1. what is the concentration of the drug in (w/v)%? 2.What is the concentration of the drug in ppm?

OpenStudy (anonymous):

1) 15 mg=0.15 g %age of drug = 0.15/6 X 100 = 2.5 % w/v 2) ppm means mg/L ppm= 15 mg/ 6 L = 2.5 ppm